OpenStudy (anonymous):

what is the area of right triangle ABC if AB=13 and CA =21

OpenStudy (anonymous):

find bc 21^2=13^2+bc^2 then area =1/2(bc)(ab) solve it ur self now

OpenStudy (anonymous):

I DONT GET IT

OpenStudy (anonymous):

come to twiddla http://www.twiddla.com/584227